OpenStudy (anonymous):

What is the slope of the line that is perpendicular to the line whose equation is 2x - y = 7? -2 -1/2 1/2 2

OpenStudy (anonymous):

2?

jimthompson5910 (jim_thompson5910):

Start by solving for y What do you get when you do so?

jimthompson5910 (jim_thompson5910):

No, 2 is not the slope of the perpendicular line

jimthompson5910 (jim_thompson5910):

But....2 is the slope of 2x-y = 7, here's why 2x - y = 7 -y = 7 - 2x -y = -2x + 7 y = -2x/(-1) + 7/(-1) y = 2x - 7 So we can see that the slope of y = 2x - 7 is 2

OpenStudy (anonymous):

Note that the equation of a perpendicular line has a slope which is the negative reciprocal of the original line.

jimthompson5910 (jim_thompson5910):

Once you have the slope of 2x - y = 7, which is 2, you then flip the fraction to get 1/2. Afterwards, you change the sign to get -1/2 So the slope of the perpendicular line is -1/2

OpenStudy (anonymous):

thank you guys